on a recent vacation, your uncle spent a total of $680 on airfare,hotel,and a rental car. the airfare cost twice as much as hotel, and the rental car was one-third as much as the hotel. use a linear system and Cramer's rule to find how much your uncle paid for each service
Answer by rfer(16322)

af=2x=408
h=x=208
rc=1/3x=68
2x+x+1/3x=680
3 1/3x=680
x=680/3 1/3
x=680*3/10
x=2040/10
x=204
2x=408
1/3x=68